A coin collector has 19 dimes and nickels with a total face value of $1.55. (They are actually worth a lot more.) How many of each coin does she have?
Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52814)

.

The number of dimes x

The number of nickels (19-x)


The total cents equation is

    10x + 5(19-x) = 155  ents,  or 


     2x + (19-x) = 31

     2x - x = 31 - 19

      x     = 12.


ANSWER.  12 dimes  and  19-12 = 7 nickels.


CHECK.   12*10 + 7*5 = 155  cents.   ! Correct !

Solved.

--------------

On coin problems,  see the lessons
    - Coin problems
    - More Coin problems
    - Solving coin problems without using equations
    - Kevin and Randy Muise have a jar containing coins
    - Typical coin problems from the archive
    - Three methods for solving standard (typical) coin word problems
    - More complicated coin problems
    - Solving coin problems mentally by grouping without using equations
    - Non-typical coin problems
    - Santa Claus helps solving coin problem
    - OVERVIEW of lessons on coin word problems
in this site.

You will find there the lessons for all levels - from introductory to advanced,
and for all methods used - from one equation to two equations and even without equations.

A convenient place to quickly observe these lessons from a  "bird flight height"  (a top view)  is the last lesson in the list.

Read them and become an expert in solution of coin problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Coin problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.




Answer by greenestamps(13203)



A student just learning algebra should certainly know how to solve the problem using formal algebra, as shown by the other tutor.

But solving a problem like this informally, with logical reasoning and simple arithmetic, is good mental exercise. It might go something like this:

(1) If all 19 coins were nickels, the total face value would be 95 cents; the actual value is $1.55, or 155 cents. The difference is 60 cents.
(2) Exchanging a nickel for a dime keeps the number of coins the same and increases the face value by 5 cents.
(3) The number of times a nickel needs to be exchanged for a dime, to make up the difference of 60 cents, is 60/5 = 12.
(4) Therefore, the collection contains 12 dimes and 19-12 = 7 nickels.